OpenStudy (anonymous):

If Php 56,734 is the present value of P hp 254,652 when will it earn its interest if money is worth 10.25% compound quartely ?

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

@jacalneaila I thought you're doing statics, suddenly you're in finite math? lol

OpenStudy (anonymous):

im done in statics :)

OpenStudy (anonymous):

thank you very much :)

OpenStudy (mathmate):

Well, this is an application of the compound interest formula. Are you familiar with the compound interest formula?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

can you check my answer :)

OpenStudy (mathmate):

sure, what do you get, and how did you get it?

OpenStudy (anonymous):

F=P(\[F=P(1+r/m)^t^m\]

OpenStudy (mathmate):

Good, what you you use for r and m?

OpenStudy (anonymous):

\[254,652=56734(1+10.25%/4)^(4)(x) \]

OpenStudy (anonymous):

m=4

OpenStudy (anonymous):

r=10.25%

OpenStudy (anonymous):

date is missing

OpenStudy (anonymous):

the answer is 14.84

OpenStudy (mathmate):

Can you continue the formula from here? 254,652=56734(1+10.25

OpenStudy (anonymous):

\[254,652=56734(1+10.25%/4)^(x)(4)\]

OpenStudy (anonymous):

254,652=56734(1+10.25%/4)^(x)(4)

OpenStudy (anonymous):

@mathmate 254,652=56734(1+10.25%/4)^(x)(4)

OpenStudy (mathmate):

Try not to write the %, and the brackets at the end are not correct. It should look like this. 254,652=56734(1+0.1025/4)^(4x) How did you solve for x?

OpenStudy (anonymous):

14.84

OpenStudy (anonymous):

i put in calcu , shift solve

OpenStudy (mathmate):

What kind of calculator solves this problem the way you entered it?

OpenStudy (anonymous):

i use 991es calcu

OpenStudy (mathmate):

Amazing, did it give more than two places of decimals? What's the rest of the number, if you have it?

OpenStudy (anonymous):

yeass

OpenStudy (mathmate):

Just curious, what's the rest of the numbers? (i.e. more accuracy for the answer?)

OpenStudy (anonymous):

just put a x in calcu

OpenStudy (anonymous):

althought , you have to adjust the decimal numbers

OpenStudy (mathmate):

The problem is, 14.84 is not the correct answer, because you will see that if you use the same compound interest formula, and solve for FV, you get PhP 254757.28 and not 254652. You need more accuracy.

OpenStudy (anonymous):

the questions is when ? it means a date

OpenStudy (mathmate):

That's ok if the answer satisfies your teacher!

OpenStudy (anonymous):

i have also a problem with this questions

OpenStudy (anonymous):

At what rate compounded monthly will php 36000 become php 48500 at the end of 85 months

OpenStudy (anonymous):

in calcu , cant solve

OpenStudy (anonymous):

@mathmate i use same formula , but icant solve

OpenStudy (mathmate):

Can you post the formula the way you put it in and I can see why, or what we can do to make it solve?

OpenStudy (mathmate):

The reason is you are not solving for the same unknown, you need some adjustments to help the calculator.

OpenStudy (anonymous):

48500=36000(1+X/12)^(12)(85/12) rate is the missing

OpenStudy (mathmate):

We will start by understanding how compound interest works, then we will be able to solve for anything, even when the 991es gets stuck, ok?

OpenStudy (anonymous):

ok :)

OpenStudy (mathmate):

So the compound interest formula works like this: F=P(1+i/m)^(mt) where F=future value, P=present value, m=number of times it compounds in a year, n=number of years, and i=(nominal) interest rate. I think you are familiar with it, right?

OpenStudy (anonymous):

48500=36000(1+x/12)^12 (85/12)?

OpenStudy (mathmate):

Not too fast, we need to understand, and modify the formula to help the calculator. I was actually surprised that it found the answer the first time around!

OpenStudy (mathmate):

Your formula is correct, but the problem is too hard for the calculator!

OpenStudy (mathmate):

To continue our discussion, are you familiar with: F=P(1+i/m)^(mt) ?

OpenStudy (anonymous):

okay okay :) no

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah

OpenStudy (mathmate):

But that's the same formula as the first AND second problem that you used!

OpenStudy (mathmate):

ok, so you're familiar with it!

OpenStudy (anonymous):

i will use it , wait :)

OpenStudy (mathmate):

Are you able to solve for the period, using the formula, and not using the automatic solver of the calculator?

OpenStudy (anonymous):

i will try :)

OpenStudy (mathmate):

The key is to take log on both sides, which then reduces the exponentiation to a multiplication. See if you can come up with a new formula using logs!

OpenStudy (mathmate):

without using the calculator, because we are not working with numbers!

OpenStudy (anonymous):

:( , i did not get it

OpenStudy (mathmate):

You're still using the calculator. I need you to take log on both sides of the formula, and simplify the equation. It will be in a form that the calculator will solve!

OpenStudy (mathmate):

For now, ditch the calculator, and use pencil and paper!

OpenStudy (anonymous):

how can i solve it

OpenStudy (anonymous):

i have a pencil and paper

OpenStudy (anonymous):

first step

OpenStudy (anonymous):

how can u teach m

OpenStudy (anonymous):

:(

OpenStudy (mathmate):

Take log on both sides, and change the exponent (on the right of the equal sign) to multiplication, as the laws of exponents require. log(ab)=log(a)+log(b) log(a^b)= b log(a) and put the modified formula for the calculator to solve. Tell me what you get.

OpenStudy (mathmate):

I got between 4 and 4.5 %.

OpenStudy (anonymous):

i got 4.1254759065 x 10^-4

OpenStudy (mathmate):

You seem to be a factor of approximately 10 off somewhere. Use this interest rate and the compound interest formula to see you you get back 48500 after 85 months. Can you show me what you put in the calculator, perhaps a 12 is missing or misplaced somewhere!

OpenStudy (mathmate):

Using interest rate of 4.1254759065 x 10^-4, we get only 36105 after 85 monthes.

OpenStudy (mathmate):

Try your calculator with log(F/P)=log((1+x/12)^(12*85/12)